Grade 12th passPhysical Chemistry

A 500 mL sample of the effluent from a water softener required six drops of standard soap solution to produce a permanent lather. The soap solution had been calibrated against an artificial hard water solution containing 0.136 g of CaCl2 per liter. On the average it required 28 drops of standard soap solution to lather 500 mL of the aritificial solution. Calculate the hardness of the effulent sample expressed as ppm of CaCO3. The hardness of the effulent sample is clearly 6/28 of that of the CaCl2 solution. It remains to express this in the standard units, assuming that each mole of CaCl2 is equivalent to a mole of CaCO3 in natural hard water?

Mehreen Jabeen
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
Hardness of effluent sample is 6/28 th of hardness of CaCl2 solution
111 parts of CaCl2 = 100 parts of CaCO3
0.136 parts (mol. wt.) of CaCl2 = 100 * 0.136 / 111 = 0.123 parts of CaCO3
Hence hardness of CaCl2 = 123 ppm
Hence hardness of effluent sample = 123 * 6 / 28 = 26 ppm
 
Regards
Arun (askIITians forum expert)